OpenStudy (anonymous):

(2x^2tan(x))/sec(x) find f ' (x) and f ' (3)

OpenStudy (anonymous):

Answer is 2 x (x cos(x)+2 sin(x)), I just don't know how to do it

OpenStudy (anonymous):

okay I got cha

OpenStudy (3psilon):

it gets a little messy

OpenStudy (anonymous):

What do you do after you get the numerator?

OpenStudy (3psilon):

Actually forget what I said. Just use the quotient rule

OpenStudy (anonymous):

(4x)(tan(x))-(2x^2)(sec(x^2)) derivative of tan(x) = sec(x)^2

OpenStudy (anonymous):

is that right for the numerator?

OpenStudy (3psilon):

\[\frac{ \sec(x)(2x^{2}\sec^{2}x+4xtanx)- 2x^{2} \tan(x)\tan(x)\sec(x) }{ \sec^{2}x }\]

OpenStudy (3psilon):

That's when it is in quotient rule form. As I said . it does get messy :/

OpenStudy (anonymous):

Wow... sorry it took long to reply, the website isn't working well for me.

OpenStudy (anonymous):

m trying to follow what you got and how you got it with the formula, i'm just learning the quotient rule

OpenStudy (3psilon):

The quotient rule is low d(high) - high d(low) all over Low low or low^2

OpenStudy (3psilon):

high is the numerator and low is the denominator when there is a d next to it that means the derivative

OpenStudy (3psilon):

In this case the d(high) involved the product rule that's why it got so messy

OpenStudy (anonymous):

okay im following now. Goodness this hard on the head... so first we found out what the numerator is from the product rule, and then we use the quotient rule

OpenStudy (3psilon):

Just take it one step at a time :) Use the quotient rule for the whole thing. But you'll come across product rule inside it